# Decentralized Zeno-Free Event-Triggered Control For Multiple Networks Subject to Stochastic Network Delays and Poisson Pulsing Attacks

### v1

@@ IV Simulations
for $\tilde{\varrho}_{i}(x_{p,i}):=0\cdot|e_{i}|$, since $-\kappa_{i}h_{i}\tilde{\Theta}^{T}_{i}\tilde{\epsilon}_{i}\leq\frac{1}{2}\kappa_{i}(\tilde{\Theta}^{2}_{i}+\tilde{\epsilon}^{2}_{i})$ holds following Yong’s inequality, where
@@
$\begin{split}&-\frac{9}{10}(\kappa_{i}\tilde{\epsilon}^{T}_{i}\tilde{\epsilon}_{i}+\lambda_{\min}(K_{i})\tilde{\Theta}^{T}_{i}\tilde{\Theta}_{i})+\frac{1}{2}\kappa_{i}(\tilde{\Theta}^{2}_{i}+\tilde{\epsilon}^{2}_{i})\\ =&-(0.9-0.5)\kappa_{i}|\tilde{\epsilon}_{i}|^{2}-(0.9\lambda_{\min}(K_{i})-0.5\kappa_{i})|\tilde{\Theta}_{i}|^{2}\\ =&-0.6|\tilde{\epsilon}_{i}|^{2}-1.05|\tilde{\Theta}_{i}|^{2}\\ \leq&-0.5525|\tilde{\epsilon}_{i}|^{2}-0.17|\tilde{\Theta}_{i}|^{2}\\ =&-((\frac{1}{2}|h_{i}|+J^{-1}_{i}\kappa_{i})^{2}+(\frac{1}{2}|h_{i}|+J^{-1}_{i}\kappa_{i})\ast|J^{-1}_{i}K_{i}|)|\tilde{\epsilon}_{i}|^{2}\\ &+(|J^{-1}_{i}K_{i}|^{2}+(\frac{1}{2}|h_{i}|+J^{-1}_{i}\kappa_{i})\ast|J^{-1}_{i}K_{i}|)|\tilde{\Theta}_{i}|^{2}\\ \leq&-((\frac{1}{2}|h_{i}|+J^{-1}_{i}\kappa_{i})|\tilde{\epsilon}_{i}|+|J^{-1}_{i}K_{i}|\tilde{\Theta}_{i}|)^{2}=-H^{2}_{l_{l},i},\end{split}$

### v3

@@ IV Simulations
where $H_{l_{i},i}:=(\frac{1}{2}|h_{i}|+J^{-1}_{i}\kappa_{i})|\tilde{\epsilon}_{i}|+|J^{-1}_{i}K_{i}\tilde{\Theta}_{i}|=0.65|\tilde{\epsilon}_{i}|+0.2|\tilde{\Theta}_{i}|$ for $\tilde{\varrho}_{i}(x_{p,i}):=0\cdot|e_{i}|$, since $-\kappa_{i}h_{i}\tilde{\Theta}^{T}_{i}\tilde{\epsilon}_{i}\leq\frac{1}{2}\kappa_{i}(\tilde{\Theta}^{2}_{i}+\tilde{\epsilon}^{2}_{i})$ holds following Yong’s inequality, where
@@
$\begin{split}&-\frac{9}{10}(\kappa_{i}\tilde{\epsilon}^{T}_{i}\tilde{\epsilon}_{i}+\lambda_{\min}(K_{i})\tilde{\Theta}^{T}_{i}\tilde{\Theta}_{i})+\frac{1}{2}\kappa_{i}(\tilde{\Theta}^{2}_{i}+\tilde{\epsilon}^{2}_{i})\\ =&-(0.9-0.5)\kappa_{i}|\tilde{\epsilon}_{i}|^{2}-(0.9\lambda_{\min}(K_{i})-0.5\kappa_{i})|\tilde{\Theta}_{i}|^{2}\\ =&-0.6|\tilde{\epsilon}_{i}|^{2}-1.05|\tilde{\Theta}_{i}|^{2}\\ \leq&-0.5525|\tilde{\epsilon}_{i}|^{2}-0.17|\tilde{\Theta}_{i}|^{2}\\ =&-((\frac{1}{2}|h_{i}|+J^{-1}_{i}\kappa_{i})^{2}+(\frac{1}{2}|h_{i}|+J^{-1}_{i}\kappa_{i})\ast|J^{-1}_{i}K_{i}|)|\tilde{\epsilon}_{i}|^{2}\\ &-(|J^{-1}_{i}K_{i}|^{2}+(\frac{1}{2}|h_{i}|+J^{-1}_{i}\kappa_{i})\ast|J^{-1}_{i}K_{i}|)|\tilde{\Theta}_{i}|^{2}\\ \leq&-((\frac{1}{2}|h_{i}|+J^{-1}_{i}\kappa_{i})|\tilde{\epsilon}_{i}|+|J^{-1}_{i}K_{i}|\tilde{\Theta}_{i}|)^{2}=-H^{2}_{l_{l},i},\end{split}$